OpenStudy (anonymous):

Medals shall be rewarded. Thanks for the help (Trigonometry question) An automobile travels at 100km/h. a) Find the angular velocity of a tire with radius 36cm. b) Through what angle will the tire turn in 30s at this speed?

OpenStudy (anonymous):

100km/h =100,000m/h =1,000/36m/s=27.8m/s assuming the tires are spinning without slipping, omega*R=v implies that 27.8/R=omega (omega is angular velocity)

OpenStudy (anonymous):

with zero acceleration theta=omega(t) where theta is the angle described and t is in seconds

OpenStudy (anonymous):

(thats omega*t not omega as a function of t)

OpenStudy (anonymous):

According to the book I have, the answer is 77.16

OpenStudy (anonymous):

what units is that in, and is that the answer to a or b

OpenStudy (anonymous):

watch out, your radius is given in centimeters but the equations require it to be in meters. (these physics questions do that a lot). The equations i wrote above should look familiar.